OpenStudy (anonymous):

medal and ill be your fan if you give me the correct answer The temperature was 68°F. Later that day the temperature was 82°F. How much does the temperature need to rise or fall to return to the original temperature of 68°F? A) Fall 14°F B) Fall 16°F C) Rise 14°F D) Rise 16°F

OpenStudy (anonymous):

a

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

68+x=82 -68 -68 x=14 so fall 14 is answer

OpenStudy (anonymous):

I just wanted to show you how to do it.

OpenStudy (anonymous):

ok thankyou so much for that

OpenStudy (anonymous):

No Problem.